# Understanding Prompt Tuning and In-Context Learning via Meta-Learning

### v1

@@ Appendix F List of Notation
$N_{\text{train}}$ Pretraining sequence length $=100$.
$N_{\text{tune}}$ Tuning sequence length $=50$.
$N_{\text{eval}}$ Evaluation sequence length $=200$.
@@
$L$ Prefix length $=6$ (or $25$ for control experiments).
$\mathcal{S}$ ‘Alphabet’ for prefix. Depends on tuning method.
@@
$\pi_{\theta}$ (Neural) parametric sequential predictor. Function from :$:\{\mathbb{R}^{D}\}^{*}\rightarrow\Delta\mathcal{A}$.
@@
$=\hat{\theta}$ if prefix-tuning and net is pretrained optimally,
$=$ at random initialization if prefix-tuning and net is untrained,
$=$ tuned weights if weight-tuning.
@@ G.1 Petraining on Random Coins, fine-tune to Single Coin
Full details on the regret curves and tuning loss curves are given in Fig. A5, and the impact of tuned prefixes on the networks’ internal dynamics is shown in Fig. 2. To interpret the latter, see Fig. A7, which shows that the internal state of models pretrained on Random Coins is highly structured, with one of the two principal components corresponding to the step $n$ and the other to the heads-to-tails ratio.
@@ G.2 Pretraining on Random Coins, fine-tune to Two-Coin Mixture
Full details on the regret curves and tuning loss curves are given in Fig. A6, and the impact of tuned prefixes on the networks’ internal dynamics is shown in Fig. A7.
@@
Figure A7: Top: Internal state (2D PCA projection) of pretrained models is highly structured—one component tracks the step $n$ and the other component tracks heads-to-tails ratio. Middle and bottom: Illustration of how different tuned prefixes manipulate the pretrained Transformer’s (middle) and LSTM’s (bottom) internal state and affect subsequent dynamics. Colored lines are from target distribution (Two-Coin Mixture), gray lines are from pretraining distribution (uniform random coins), same setting as in Fig. A6. Red circles mark the end of the prefixes.
@@ G.3 Untrained network, fine-tuned on Two-Coin Mixture
Full details on the regret curves and tuning loss curves are given in Fig. A8.
@@ G.4 Untrained network, fine-tuned on Random Coins
Full details on the regret curves and tuning loss curves are given in Fig. A9.
@@
Our main experiments use relatively short prefixes of length $L=6$. The main point is to demonstrate how powerful even very short soft prefixes can be. Additionally, the number of possible hard prefixes grows exponentially with $L$, making exhaustive hard token search for long prefixes intractable. To confirm that our negative result on tuning the pretrained predictor to a Two-Coin Mixture Fig. 3 is not an artifact of insufficient prefix length, we repeat the soft prefix-tuning experiments with more than triple the prefix length of $L=25$. Results are shown in Fig. A10. We also show in the same figure that increasing the soft prefix length for tuning untrained networks has only marginal effect.
@@
Figure A10: prefix-tuning to Two-Coin Mixture with prefix length $L=25$ (in contrast to all other experiments where $L=6$). Compare pretrained results with Fig. 3, and untrained results with Fig. 4. Despite a more than tripling the prefix length, only marginal increases in ‘SoftPF’ performance can be seen in some cases. It is not enough to reach Byaes-optimality on the target distribution, which means that our qualitative results hold.

### v2

@@ Appendix F List of Notation
$N_{\text{train}}$ Pretraining sequence length =100=100.
$N_{\text{tune}}$ Tuning sequence length =50=50.
$N_{\text{eval}}$ Evaluation sequence length =200=200.
@@
$L$ Prefix length =6=6 (or $25$ for control experiments).
$\mathcal{S}$ “Alphabet” for prefix. Depends on tuning method.
@@
$\pi_{\theta}$ (Neural) parametric sequential predictor. Function from $\{\mathbb{R}^{D}\}^{*}\rightarrow\Delta\mathcal{A}$.
@@
$\hat{\theta}$ if prefix-tuning and net is pretrained optimally,
== at random initialization if prefix-tuning and net is untrained,
== tuned weights if weight-tuning.
@@ G.1 Petraining on Random Coins, fine-tune to Single Coin
Full details on the regret curves and tuning loss curves are given in Fig.˜A5, and the impact of tuned prefixes on the networks’ internal dynamics is shown in Fig.˜2. To interpret the latter, see Fig.˜A7, which shows that the internal state of models pretrained on Random Coins is highly structured, with one of the two principal components corresponding to the step $n$ and the other to the heads-to-tails ratio.
@@ G.2 Pretraining on Random Coins, fine-tune to Two-Coin Mixture
Full details on the regret curves and tuning loss curves are given in Fig.˜A6, and the impact of tuned prefixes on the networks’ internal dynamics is shown in Fig.˜A7.
@@
Figure A7: Top: Internal state (2D PCA projection) of pretrained models is highly structured—one component tracks the step $n$ and the other component tracks heads-to-tails ratio. Middle and bottom: Illustration of how different tuned prefixes manipulate the pretrained Transformer’s (middle) and LSTM’s (bottom) internal state and affect subsequent dynamics. Colored lines are from target distribution (Two-Coin Mixture), gray lines are from pretraining distribution (uniform random coins), same setting as in Fig.˜A6. Red circles mark the end of the prefixes.
@@ G.3 Untrained network, fine-tuned on Two-Coin Mixture
Full details on the regret curves and tuning loss curves are given in Fig.˜A8.
@@ G.4 Untrained network, fine-tuned on Random Coins
Full details on the regret curves and tuning loss curves are given in Fig.˜A9.
@@
Our main experiments use relatively short prefixes of length $L=6$. The main point is to demonstrate how powerful even very short soft prefixes can be. Additionally, the number of possible hard prefixes grows exponentially with $L$, making exhaustive hard token search for long prefixes intractable. To confirm that our negative result on tuning the pretrained predictor to a Two-Coin Mixture in Fig.˜3 is not an artifact of insufficient prefix length, we repeat the soft prefix-tuning experiments with more than triple the prefix length of $L=25$. Results are shown in Fig.˜A10. We also show in the same figure that increasing the soft prefix length for tuning untrained networks has only marginal effect.
@@
Figure A10: prefix-tuning to Two-Coin Mixture with prefix length $L=25$ (in contrast to all other experiments where $L=6$). Compare pretrained results with Fig.˜3, and untrained results with Fig.˜4. Despite a more than tripling the prefix length, only marginal increases in ‘SoftPF’ performance can be seen in some cases. It is not enough to reach Byaes-optimality on the target distribution, which means that our qualitative results hold.
Appendix I Reducing the embedding dimensionality
Results shown in Fig.˜A11 reveal that the superiority of ‘SoftPT’ over the other soft prefix tuning methods is largely explained by the much higher dimensionality of the embedding vectors ($128$-dimensional in the main experiments; now reduced to $4$ dimensions), compared to input vectors (which are two-dimensional). In frontier models, the input dimensionality is typically higher than the embedding dimensionality, which could in principle result in soft input prefix tuning methods like ‘RealPT’ outperforming embedding tuning. We leave the question of which prefix tuning method works best at frontier model scale to the large and active research community. Note that since the “width” of our Transformers is equal to the embedding dimensionality (except the width of the MLP layer inside the attention block), the Transformer in our reduced embedding dimensionality experiments is much smaller compared to the main experiments, whereas the LSTM has the same size after the embedding layer.
(a) Pretrained Transformer to Single Coin.
(b) Pretrained LSTM to Single Coin.
(c) Pretrained Transformer to Two-Coin Mixture.
(d) Pretrained LSTM to Two-Coin Mixture.
Figure A11: Reducing the embedding dimensionality to 44 (compared to $128$ in main experiments) largely eliminates the superiority of ‘SoftPT’ compared to ‘RealPT’ (and shortens the gap to the other prefix tuning methods), revealing that the superior performance observed in the main experiments is largely explained by the many more degrees of freedom when tuning soft embedding prefixes vs. soft input prefixes. Compare the results shown here (particularly the dark blue ‘SoftPT’ bar) against Fig.˜1 and Fig.˜4 in the main paper. Plots show median results over 33 repetitions.
Appendix J Experiments with larger networks
Fig.˜A12 shows results for increasing the network size. Compared to the main experiments we double the embedding dimensionality ($128\rightarrow 256$), the width of layers ($128\rightarrow 256$), and the number of layers ($1\rightarrow 2$). Qualitatively, our main claims hold. Particularly, that prefix tuning can be used to optimally adapt the pretrained predictor to the Single Coin target distribution, but cannot be used for perfect adaptation to the Two-Coin Mixture task. Anecdotally, we have observed our main results to hold robustly, as long as the network size and number of training and tuning steps is large enough. For too small networks, or networks trained or tuned too little, results become more inconsistent. From a theoretical perspective, too small networks violate the realizability condition, and networks with too little training violate the convergence condition.
(a) Pretrained Transformer to Single Coin.
(b) Pretrained LSTM to Single Coin.
(c) Pretrained Transformer to Two-Coin Mixture.
(d) Pretrained LSTM to Two-Coin Mixture.
Figure A12: Results for larger networks compared to main experiments are qualitatively equivalent, and show that our main findings are robust against changing model size. Compare the results shown here (particularly the dark blue ‘SoftPT’ bar) against Fig.˜1 and Fig.˜4 in the main paper. Plots show median results over 33 repetitions.